OpenStudy (anonymous):

At a large university, the probability that a student takes calculus and is on the dean's list is 0.042. the probability that the student is on the deans list is 0.21. Find the probability that the student is taking calculus, given that he or she is on the deans list.

OpenStudy (anonymous):

conditional probability - i'll have to revise it!!

OpenStudy (anonymous):

right required proibability = P(Calculus and on deans list) / P(denos list) = 0.042 / 0.21 = 0.2 the other guy was right

OpenStudy (anonymous):

so do i multiply or divide

OpenStudy (anonymous):

i dont want to give you are wrong answer Vive so I'll recheck

OpenStudy (anonymous):

OK the propabilty of C given A is P(A intersect C) / P(A) here the the top of fraction = 0.042 and P(A) = 0.21 so its divide P = 0.042 / 0.21 = 0 .2 or 20%

OpenStudy (anonymous):

here P(A) = probabilty hes on deans list and P(C) is P he takes calculas I'm confident of that

OpenStudy (anonymous):

oh ok ok making sense

OpenStudy (anonymous):

good